OpenStudy (anonymous):

Can someone help me PLEASE ! :) 10x-4(x-2)=9x+16-2x

OpenStudy (anonymous):

Are you looking for x?

OpenStudy (anonymous):

10x-4x+8=9x+16-2x 6x+8=7x+16 x=?

OpenStudy (anonymous):

itd be -8 ...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks!